OpenStudy (anonymous):

Which is the exact area of the circle? Radius of 9 feet. I got 81 but 81 pi squared, 81 pi

OpenStudy (lgbasallote):

area of a circle is pi (radius)^2 right? then area = pi (9)^2 area = pi(81)

OpenStudy (campbell_st):

thats correct

OpenStudy (anonymous):

Wait 81 sq feet or 81 pi symbol

OpenStudy (lgbasallote):

81 pi sq ft